OpenStudy (anonymous):

A child tosses five quarters in the air. Each quarter lands either falls so that “heads” is visible or “tails” is visible. What is the probability that more than one quarter lands on tails?

OpenStudy (anonymous):

@satellite73 I re-posted it. You gave 1-(1/2)^5 x 5-(1/2)^5. I am not sure how the 1 and 5 are fitting into this puzzle here.

OpenStudy (anonymous):

@mathmate @zepdrix @kropot72 if anyone else wants to walk me through this process that would be awesome.

OpenStudy (anonymous):

the probability that all land heads is \[(\frac{1}{2})^5\] or \[\frac{1}{2^5}\]is that clear?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I just don't understand how the values 1 and 5 are fitting into this

OpenStudy (anonymous):

so how about the probability that exactly one tail and four heads?

OpenStudy (anonymous):

any particular sequence has probability \[\frac{1}{2^5}\]

OpenStudy (anonymous):

there are five such sequences with 1 tail and two heads t h h h h h t h h h h h t h h h h h t h h h h h t

OpenStudy (anonymous):

each has the same probability and they are all different, so you add them up i.e. multiply by 5 to get \[\frac{5}{2^5}\]

OpenStudy (anonymous):

Okay. It was actually - 5(1/2)^5 at the end (just a quick correction).

OpenStudy (anonymous):

ok now lets go slow

OpenStudy (anonymous):

you want "more than one quarter lands on tails?"

OpenStudy (anonymous):

more than one means 2 or 3, or 4, or 5 you can compute each of those and add to get your answer, but there is a shorter way

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

but it is shorter to compute the probability you get 0 or 1

OpenStudy (anonymous):

that is what is left so compute those two, add them together, subtract the result from one that way you only have to compute two probabilities instead of 4

OpenStudy (anonymous):

i am going to try just rereading over this to see if i can't get it without wasting your time

OpenStudy (anonymous):

ok ask if you have any questions i can give another example quickly if you likke

OpenStudy (anonymous):

heh

OpenStudy (anonymous):

im just terrible at math, dude

OpenStudy (anonymous):

I can see that 5(.03125)^5 is each combination where there is 1 heads

OpenStudy (anonymous):

but i dont understand the rest of it.

OpenStudy (kropot72):

The probability of zero tails on all 5 coins is (1/2)^5. Using the binomial distribution, the probability of exactly one tail among the 5 coins is given by: \[\large P(1\ tail)=5C1\times(\frac{1}{2})^{1}\times(\frac{1}{2})^{4}=\frac{5}{2^{5}}\] Adding the probabilities of zero tails and one tail we get: \[\large P(0\ or\ 1\ tail)=\frac{6}{2^{5}}\] So the probability of more than one tail is \[\large P(more\ than\ 1\ tail)=1-\frac{6}{2^{5}}=you\ can\ calculate\]

zepdrix (zepdrix):

:o

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

Make sense of all this yet? :)

OpenStudy (anonymous):

no man. this last guy didn't even go 5(1/2)^5

zepdrix (zepdrix):

Hmm I'm not sure if my explanation is going to be any better, I would have done it very similar to Kropot's way.

OpenStudy (anonymous):

i feel so dumb

zepdrix (zepdrix):

We have 5 slots, two options for each slot. So `2^5 total` possible combinations of flips. How many ways are there to flip `0 tails`? 1 way, all heads. how many ways are there to flip `1 tails`? 5 different ways. So we can flip 1 tail or less in 6 ways. Then we can flip more than 1 tail in `Total` - `6` ways. These are the total number of flips which result in more than 1 tails: \(\large\rm 2^5-6\) But they want the probability, so we need to divide this amount by the total number of possible flips, our sample space. \(\large\rm \dfrac{2^5-6}{2^5}\) Doing some division gives us: \(\large\rm 1-\frac{6}{2^5}\)

zepdrix (zepdrix):

Do you have answer key for this problem? :U

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how did you pull 1 out of that :X?

zepdrix (zepdrix):

\[\large\rm \frac{2^5-6}{2^5}\quad=\quad \frac{2^5}{2^5}-\frac{6}{2^5}\]

OpenStudy (anonymous):

oh, duh, sorry.

OpenStudy (anonymous):

oh, your explanation is better. i can smell the synapses firing!

OpenStudy (anonymous):

or burning. not sure

zepdrix (zepdrix):

lol XD

zepdrix (zepdrix):

I find probability stuff to be incredibly confusing. It never made sense to me like other math does. So I feel yer pain :U

OpenStudy (anonymous):

i just have a hard time understanding what is going on inside the formulas with this

OpenStudy (anonymous):

ok, so i'm going to try to figure out what it would be for 3 tails.

zepdrix (zepdrix):

For exactly 3 tails?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how would i do that without just counting everything up? :(

zepdrix (zepdrix):

I think it's just \(\left(\begin{matrix}5 \\ 3\end{matrix}\right)\) Thinking...

OpenStudy (anonymous):

what is that c(5,3)?

zepdrix (zepdrix):

Nah that can't be right :) sec thinking XD

zepdrix (zepdrix):

yes, combination

zepdrix (zepdrix):

Answer key? XD

OpenStudy (anonymous):

that's a problem i just made up

zepdrix (zepdrix):

ah :)

zepdrix (zepdrix):

https://www.youtube.com/watch?v=udG9KhNMKJw

OpenStudy (anonymous):

i want to make sure i can use this for flipping coins when it comes up on the test tomorrow.. which i'm probably screwed on

OpenStudy (anonymous):

ok i'm watching link

zepdrix (zepdrix):

Ah yes it is simply 5 choose 3, it worked yay!

zepdrix (zepdrix):

We have 5 places to put 3 tails, order does not matter, so \(\large\rm C(5,3)=\frac{5!}{3!(5-3)!}=~...~=10\)

OpenStudy (anonymous):

well i'll be damned

OpenStudy (anonymous):

so if i wanted the probability of flipping > 2heads i could calculate it that way

zepdrix (zepdrix):

But again, if you wanted a `probability`, the odds of getting exactly 3 tails would be: \(\large\rm \dfrac{10}{2^5}\)

zepdrix (zepdrix):

Ummm

OpenStudy (anonymous):

it's not a probability?

OpenStudy (anonymous):

it's out of 1....

zepdrix (zepdrix):

Hmm? :o

zepdrix (zepdrix):

1 = 100% 10/32 = 31.25% So the odds of rolling exactly 3 tails is about 31 percent chance.

OpenStudy (anonymous):

no no no

OpenStudy (anonymous):

you've got it all wrong.

OpenStudy (anonymous):

your odds of (F)lipping 3 tails is 31%

zepdrix (zepdrix):

Ya sorry I was thinking of dice I guess :) lol rolling

OpenStudy (anonymous):

can't we just stick a p infront of it and call it a day? p(31%)

OpenStudy (anonymous):

bam!

zepdrix (zepdrix):

;p

zepdrix (zepdrix):

But to your other question, about using this shortcut for >2, Yes, that should work out fine! C(5,3) for all the combinations involving 3 tails, C(5,4) for all the combinations involving 4 tails, C(5,5) for all the combinations involving 5 tails. So all the combinations involving more than 2 tails flipped is C(5,3) + C(5,4) + C(5,5) Then the probability of flipping more than 2 tails is\[\large\rm \frac{C(5,3)+C(5,4)+C(5,5)}{2^5}\]

OpenStudy (anonymous):

yeah... did we already go over how to make this into a probability? im going to scroll up and check

zepdrix (zepdrix):

ummm

OpenStudy (anonymous):

so we kinda did just stick a p infront of it almost

zepdrix (zepdrix):

You divide your `number of possibilities in the given event` by the `total number of possibilities in the sample space` We determined earlier that the size of the space is 2^5, that's the total number of possible outcomes from flipping 5 coins. What we were doing separately was counting the number of outcomes when more than 2 tails are flipped. Then we divide that by the total for our % in relation to the total. Sorry I'm not really familiar with the P notation :\